# Ground-state phase diagram of $S=1/2$ Heisenberg model on 2D square-hexagon-octagon lattice

### v1

@@ III.2 $J_{1}$-$J_{3}$ ground-state phase diagram
Moving forward to the next step, our aim is to investigate the scenario where $\lambda_{1}=0$ and $\lambda_{2}=J_{3}/J_{1}$ varies. In this case, the system comprises an array of decoupled ladders. Therefore, to analyze this situation, we only need to focus on the $J_{1}$-$J_{3}$ Heisenberg model defined on a single ladder. Consequently, we turn to the DMRG method as a supplementary numerical approach for this particular case. Similar to the $J_{1}$-$J_{2}$ case, there are two decoupled limits corresponding to $\lambda_{2}=0$ and $\lambda_{2}=\infty$, respectively. $\lambda_{2}=0$ corresponds to decoupled hexagon limit, in the small $\lambda_{2}$ region, the system is in the hexagon phase. $\lambda_{2}=\infty$ corresponds to the ladder staggered dimer (LSD) phase which is adiabatically connected to the direct product of the isolated dimers in staggering order. Between this two phases, there may have one or more phases.
@@
Figure 4: (a) When $\lambda_{1}=0$, the Binder cumulant which is defined on a single ladder varies with $\lambda_{2}$. As the intermediate phase is nonmagnetic, the Binder cumulant fails to reveal phase transitions between nonmagnetic phases, and no intersection points are visible. (b) When $\lambda_{1}=0$, the spin stiffness in the $x$-direction changes with $\lambda_{2}$. Unlike the Binder cumulant, spin stiffness can detect phase transitions between nonmagnetic phases. However, there are four intersections, so extrapolation is needed to pinpoint the exact phase transition points. (c) and (d) show the extrapolation results for the spin stiffness intersections at the two phase transition points, with the following form of extrapolation function: $\lambda_{2,c}(L)=\lambda_{2,c}(\infty)+aL^{-\omega}$ [14]. Near $\lambda_{2,c}=1.040(2)$ and $\lambda_{2,c}=1.634(2)$, the extrapolation curves approach and intersect each other.
@@
To detect any potential phases between the hexagon phase and the LSD phase, several factors need to be considered. In a quasi-one-dimensional configuration with short-range Heisenberg interaction, quantum fluctuations are relatively strong, preventing the formation of a magnetically ordered phase in the ground state. As a result, the Binder cumulant is not suitable for characterizing phase transitions. Indeed, from the QMC results of the Binder cumulant (Fig. 4(a)), no intersections are found for different sizes. Therefore, we need to explore other physical quantities to identity the location of the phase transitions. We choose the spin stiffness along the $x$ direction to detect phase transitions, as shown in Fig. 4(b). The spin stiffness graph shows four intersections between adjacent sizes. To determine whether these four intersections each correspond to a phase transition point, we combine their extrapolation curves for comparison. It turns out that under the thermodynamic limit, two of these four points approach $\lambda_{2,c}=1.040(2)$, and the other two approach $\lambda_{2,c}=1.634(2)$. And these two points are also exactly at the peak positions of spin stiffness. Thus, there are only two phase transition points in the range of $\lambda_{1}=0$, $\lambda_{2}\in[0,4]$. These two points divide the system into three phases. However, the nature of the middle phase requires further investigation.
@@
Figure 5: (a) and (b) display the extrapolation of the excitation energy gaps for singlet ($S=0$), triplet ($S=1$) and quintuplet ($S=2$) states at $\lambda_{2}=1.4$, under periodic and open boundary condition, respectively. The fitting function used is: $\Delta(L)=a+e^{-L/\xi}(b/L+c/L^{2})$ [35, 36]. Panel (c) shows the entanglement spectrum for a linear system size $L=32$ at $\lambda_{2}=1.4$ under open boundary conditions. Pannel (d) presents the entanglement entropy near the phase transition points at $\lambda_{1}=0$, with $L=32$ under periodic boundary condition. The inset includes liner extrapolations of the central charges $c$ at the these transition points.
@@
The intermediate phase cannot be a magnetically ordered phase; instead, it may be a gapless Luttinger liquid or a gapped phase. To determine the excitation energy gaps, we use the density matrix renormalization group method to calculate the energy gaps from the singlet ground state to the excited singlet state, triplet state and quintuplet state of the finite-size system and then extrapolate to the thermodynamic limit. At $\lambda_{2}=1.4$, under periodic boundary condition (PBC), all three gaps extrapolate to finite values. This indicates that the ground state is unique and fully gapped. In contrast, under open boundary condition (OBC), the first excited triplet state becomes degenerate with ground state, While the other gaps remain finite. This suggests the system may be in a bosonic symmetry-protect topological phase with end state under OBC. To further confirm the topological nature of the intermediate phase, we present the entanglement spectrum in Fig. 5(c). The entanglement spectrum is computed from the Schmidt decomposition performed at the central bond (see dashed line in the schematic diagram of the SPT phase in Fig. 1) during the DMRG sweep. From our DMRG calculations, it is evident that all entanglement spectral levels are at least doubly degenerate which is similar to the $S=1$ Haldane chain [27]. Hence, the intermediate phase is a Haldane-like symmetry protected topological phase [26, 27].
@@
Furthermore, in this 1D system, the continuous quantum phase transitions between the Haldane-like SPT phase and the other two phases are effectively captured by the Wess-Zumino-Witten (WZW) conformal field theory [37], which provides a framework to understand the universality class of these transitions. The central charge $c$ is a key physical quantity in this context. In our DMRG calculations, we computed the entanglement entropy for varying subsystem sizes $l$ while keeping the total system size $L$ fixed. By fitting the calculated entanglement entropy data under periodic boundary condition to the Calabrese-Cardy formula $S=\frac{c}{3}ln\left(\frac{L}{\pi}sin\left(\frac{\pi l}{L}\right)\right)+constant$ (see Fig. 5 (d)) [38, 39], and then extrapolate to $L=\infty$, we are able to determine the central charges $c=1$ in the thermodynamic limit.
@@ III.3 Full ground-state phase diagram
We have already analyzed and confirmed the phases and phase transitions in the cases of $\lambda_{1}=0$ and $\lambda_{2}=0$. To complete the remaining phase diagram, we continue to apply the finite-size scaling method to extrapolate the intersection of Binder cumulants for different sizes to accurately pinpoint the phase boundaries. To demonstrate the reliability of our data, we present a detailed analysis of cases where we vary $\lambda_{2}$ while fixing $\lambda_{1}$ at 0.50.5,$1.0$, and $4$, as well as the case where we vary $\lambda_{1}$ while fixing $\lambda_{2}$ at $1.4$. The overall phase diagram for the SHO lattice is presented in Fig. 2. As can be seen, within the chosen parameter range, the SHO lattice model exhibits five distinct phases: the hexagon phase, the OSD phase, the LSD phase, the Haldane-like SPT phase, and the AFM phase. the AFM phase is surrounded by other non-magnetic phases in the phase diagram.
Figure 6: (a) and (b) show Binder cumulant $U_{2}$ as a function of $\lambda_{2}$ near the two critical points for $\lambda_{1}=0.5$. (c) shows $U_{2}$ versus $\lambda_{2}$ near the critical point for $\lambda_{1}=1$. (d) presents $U_{2}$ as a function of $\lambda_{1}$ near the critical point for $\lambda_{2}=1.4$. (e) and (f) illustrate $U_{2}$ versus $\lambda_{2}$ near the two critical points for $\lambda_{1}=4$. The insets show the extrapolations of the corresponding critical points.
In the vertical path with $\lambda_{1}$ fixed at $0.5$ and varying $\lambda_{2}$, two phase transitions occur: one from the hexagon phase to the AFM phase, and the other from the AFM phase to the ladder staggered dimer (LSD) phase. Through finite-size extrapolation, the critical points are accurately determined to be $\lambda_{2,c}=0.323(7)$ and $\lambda_{2,c}=2.040(1)$. In the vertical path with $\lambda_{1}$ fixed at $1$ and varying $\lambda_{2}$, only one quantum phase transition is observed, namely from the AFM phase to the LSD phase. The critical point for this transition is accurately identified as $\lambda_{2,c}=2.320(2)$. At the point where $\lambda_{1}=\lambda_{2}=1$, the system is topologically equivalent to the square-lattice Heisenberg model. In the thermodynamic limit, the staggered magnetic order is extrapolated to approximately $0.307$ as shown in Fig. 7, which aligns closely with previously reported results [40]. When $\lambda_{1}=1$ and $\lambda_{2}=0$, the magnetic order is reduced compared to the case where $\lambda_{1}=\lambda_{2}=1$. In the vertical path with $\lambda_{1}$ fixed at $4$ and varying $\lambda_{2}$, the AFM phase region shrinks compared to cases with smaller $\lambda_{1}$ values. Two continuous quantum phase transitions occur between the AFM phase and the orthogonal staggered dimer (OSD) phase, as well as between the AFM phase and the ladder staggered dimer (LSD) phase. The critical points for these transitions are accurately determined to be $\lambda_{2,c}=3.939(6)$ and $\lambda_{2,c}=4.228(4)$. And the AFM order at the $\lambda_{1}=\lambda_{2}=4$ point is very small compared to other AFM region, as shown in Fig. 7. To illustrate the quantum phase transition between the Haldane-like SPT phase and the AFM phase, we consider a horizontal path where $\lambda_{1}$ is varied while $\lambda_{2}$ is fixed at $1.4$. As shown in Fig. 2 (d), the Binder cumulants for different system sizes exhibit intersections that remain relatively unchanged with increasing system sizes. By performing extrapolation, we accurately determine a quantum critical point at $\lambda_{1,c}=0.106(1)$. This is analogous to the case of coupled Haldane chains transitioning into a 2D square lattice, where a similarly small critical interchain interaction has been reported [41].
We plot the identified phase transition points on the phase diagram and use the same finite-size scaling method to determine additional points across the phase diagram. This process results in a detailed and comprehensive phase diagram for the Heisenberg model on 2D SHO lattice, as displayed in Fig. 2. Due to the presence of an even sublattice, the AFM phase tends to emerge readily. It occupies a middle area of the phase diagram and is surrounded by four other non-magnetic phases.
@@
Figure 7: The extrapolation curves of the squared magnetization for different parameters settings $\lambda_{1}=\lambda_{2}=4$, $\lambda_{1}=\lambda_{2}=1$, and $\lambda_{1}=1$, $\lambda_{2}=0$. Their extrapolated values are $M^{2}_{s}(\infty)=0.0061(7),0.0944(3),0.0609(2)$ respectively.
@@ III.4 Critical behaviors
After studying the position of each phase in the model clearly, we continue to explore the critical behavior in the phase transition. Since the antiferromagnetic phase breaks the spin $SU(2)$ continuous symmetry in this model while the other phases such as dimer phase or hexagon phase do not break this continuous symmetry and translation symmetry, it is reasonable to conjecture that the phase transition of this model should belong to the 3D $O(3)$ universality class. At present, there are precise critical exponents values for $O(3)$ universality class in the literature [42]. We only need to do the data collapse of the Binder cumulant $U_{2}$ and compare the obtained critical exponent with the known value for confirmation. The finite-size scaling form of Binder cumulant is given by $U_{2}(\lambda_{1/2},L)=f[L^{1/\nu}(\lambda_{1/2}-\lambda_{1/2,c})/\lambda_{1/2,c}]$, where $f$ is scaling function, $\lambda_{1/2,c}$ is critical point, and $\nu$ is the critical exponent of correlation length. To extract the optimal critical point and critical exponent from the finite-size Binder cumulant, we employ the scaling analysis method based on Gaussian process regression proposed in Refs. [43, 44], which enables data from different system sizes to collapse onto a single smooth scaling function at the optimal best-fit critical parameters. We select representative paths from the four phase boundaries in the phase diagram and apply the above scaling analysis method to determine their critical exponents and judge all phase transition types accordingly.
Figure 8: Panels (a)-(d) show the data collapse of the Binder cumulant $U_{2}$ for the phase transitions between AFM phase and hexagon phase ($\lambda_{1}=0.5$), LSD phase ($\lambda_{1}=1.0$), OSD phase ($\lambda_{1}=2.0$), and Haldane-like SPT phase ($\lambda_{2}=1.4$), respectively. Here, $\lambda_{1/2,c}$ and $\nu$ represent the optimal critical point and critical exponent, respectively. The error estimation is based on 95% confidence interval of Gaussian distribution.
@@
The results of data collapse are shown in Fig. 8, where the best estimates of the critical point $\lambda_{1/2,c}$ and critical exponent $\nu$ for different phase transitions are presented. The critical points obtained here are all consistent with the phase transition points shown in Fig. 2. The critical exponents $\nu$ we obtain here are all in good agreement with the $O(3)$ value $\nu\approx 0.7112$ [42], indicating that the phase transition from the four non-magnetic phases to AFM phase in the Heisenberg model on SHO lattice indeed belong to the 3D $O(3)$ universality class.

### v2

@@ III.2 $J_{1}\text{-}J_{3}$ ground-state phase diagram
Moving forward to the next step, our aim is to investigate the scenario where $\lambda_{1}=0$ and $\lambda_{2}=J_{3}/J_{1}$ varies. In this case, the system comprises an array of decoupled ladders. Therefore, to analyze this situation, we only need to focus on the $J_{1}\text{-}J_{3}$ Heisenberg model defined on a single ladder. Consequently, we turn to the DMRG method as a supplementary numerical approach for this particular case. Similar to the $J_{1}\text{-}J_{2}$ case, there are two decoupled limits corresponding to $\lambda_{2}=0$ and $\infty$, respectively. $\lambda_{2}=0$ corresponds to decoupled hexagon limit, and in the small $\lambda_{2}$ region the system is in the hexagon phase. $\lambda_{2}=\infty$ corresponds to the LSD phase which is adiabatically connected to the direct product of the isolated dimers in staggering order. Between these two phases, there may be one or more phases.
@@
Figure 4: (a) When $\lambda_{1}=0$, the Binder cumulant which is defined on a single ladder varies with $\lambda_{2}$. As the intermediate phase is nonmagnetic, the Binder cumulant fails to reveal phase transitions between nonmagnetic phases, and no intersection points are visible. (b) When $\lambda_{1}=0$, the spin stiffness in the $x$ direction changes with $\lambda_{2}$. Unlike the Binder cumulant, spin stiffness can detect phase transitions between nonmagnetic phases. However, there are four intersections, so extrapolation is needed to pinpoint the exact phase transition points. (c, d) The extrapolation results for the spin stiffness intersections at the two phase transition points, with the following form of the extrapolation function: $\lambda_{2,c}(L)=\lambda_{2,c}(\infty)+aL^{-\omega}$ [14]. Near $\lambda_{2,c}=1.040(2)$ and $1.634(2)$, the extrapolation curves approach and intersect each other.
@@
To detect any potential phases between the hexagon phase and the LSD phase, several factors need to be considered. In a quasi-one-dimensional configuration with short-range Heisenberg interaction, quantum fluctuations are relatively strong, preventing the formation of a magnetically ordered phase in the ground state. As a result, the Binder cumulant is not suitable for characterizing phase transitions. Indeed, from the QMC results of the Binder cumulant [Fig. 4(a)], no intersections are found for different sizes. Therefore, we need to explore other physical quantities to identity the location of the phase transitions. We choose the spin stiffness along the $x$ direction to detect phase transitions, as shown in Fig. 4(b). The spin stiffness graph shows four intersections between adjacent sizes. To determine whether these four intersections each correspond to a phase transition point, we combine their extrapolation curves for comparison. It turns out that under the thermodynamic limit, two of these four points approach $\lambda_{2,c}=1.040(2)$, and the other two approach $\lambda_{2,c}=1.634(2)$. And these two points are also exactly at the peak positions of spin stiffness. Thus, there are only two phase transition points in the range of $\lambda_{1}=0$, $\lambda_{2}\in[0,4]$. These two points divide the system into three phases. However, the nature of the middle phase requires further investigation.
@@
Figure 5: (a, b) The extrapolation of the excitation energy gaps for singlet ($S=0$), triplet ($S=1$), and quintuplet ($S=2$) states at $\lambda_{2}=1.4$, under periodic and open boundary conditions, respectively. The fitting function used is $\Delta(L)=a+e^{-L/\xi}(b/L+c/L^{2})$ [35, 36]. (c) The entanglement spectrum for a linear system size $L=32$ at $\lambda_{2}=1.4$ under open boundary conditions. (d) The entanglement entropy near the phase transition points at $\lambda_{1}=0$, with $L=32$ under periodic boundary conditions. The inset includes linear extrapolations of the central charges $c$ at the these transition points.
@@
The intermediate phase cannot be a magnetically ordered phase; instead, it may be a gapless Luttinger liquid or a gapped phase. To determine the excitation energy gaps, we use the density matrix renormalization group method to calculate the energy gaps from the singlet ground state to the excited singlet state, triplet state, and quintuplet state of the finite-size system and then extrapolate to the thermodynamic limit. At $\lambda_{2}=1.4$, under PBCs, all three gaps extrapolate to finite values [see Fig. 5(a)]. This indicates that the ground state is unique and fully gapped. In contrast, under OBCs, the first excited triplet state becomes degenerate with the ground state, while the other gaps remain finite, as can be seen in Fig. 5(b). This suggests the system may be in a bosonic SPT phase with end state under OBCs. To further confirm the topological nature of the intermediate phase, we present the entanglement spectrum in Fig. 5(c). The entanglement spectrum is computed from the Schmidt decomposition performed at the central bond (see dashed line in the schematic diagram of the SPT phase in Fig. 1) during the DMRG sweep. From our DMRG calculations, it is evident that all entanglement spectral levels are at least doubly degenerate which is similar to the $S=1$ Haldane chain [27]. Hence, the intermediate phase is a Haldane-like SPT phase [26, 27].
@@
Furthermore, in this one-dimensional system, the continuous quantum phase transitions between the Haldane-like SPT phase and the other two phases are effectively captured by the Wess-Zumino-Witten conformal field theory [37], which provides a framework to understand the universality class of these transitions. The central charge $c$ is a key physical quantity in this context. In our DMRG calculations, we computed the entanglement entropy for different subsystem sizes $l$ while keeping the total system size $L$ fixed [see Fig. 5(d)]. By fitting the calculated entanglement entropy data under periodic boundary conditions to the Calabrese-Cardy formula $S=\frac{c}{3}ln\left(\frac{L}{\pi}\mathrm{sin}\left(\frac{\pi l}{L}\right)\right)+\mathrm{const}$ [see Fig. 5 (d)] [38, 39], and then extrapolating to $L=\infty$, we are able to determine the central charges $c=1$ in the thermodynamic limit.
Figure 6: (a, b) Binder cumulant $U_{2}$ as a function of $\lambda_{2}$ near the two critical points for $\lambda_{1}=0.5$. (c) $U_{2}$ vs $\lambda_{2}$ near the critical point for $\lambda_{1}=1$. (d) $U_{2}$ as a function of $\lambda_{1}$ near the critical point for $\lambda_{2}=1.4$. (e, f) $U_{2}$ vs $\lambda_{2}$ near the two critical points for $\lambda_{1}=4$. The insets show the extrapolations of the corresponding critical points.
@@ III.3 Full ground-state phase diagram
We have already analyzed and confirmed the phases and phase transitions in the cases of $\lambda_{1}=0$ and $\lambda_{2}=0$. To complete the remaining phase diagram, we continue to apply the finite-size scaling method to extrapolate the intersection of Binder cumulants for different sizes to accurately pinpoint the phase boundaries. To demonstrate the reliability of our data, we present a detailed analysis of cases where we vary $\lambda_{2}$ while fixing $\lambda_{1}$ at 0.50.5,$1.0$, and $4$, as well as the case where we vary $\lambda_{1}$ while fixing $\lambda_{2}$ at $1.4$, as presented in Fig. 6. The overall phase diagram for the SHO lattice is presented in Fig. 2. Within the chosen parameter range, the SHO lattice model exhibits five distinct phases: the hexagon phase, the OSD phase, the LSD phase, the Haldane-like SPT phase, and the AFM phase. The AFM phase is surrounded by other nonmagnetic phases in the phase diagram.
@@
Figure 7: The extrapolation curves of the squared magnetization for different parameter settings $\lambda_{1}=\lambda_{2}=4$; $\lambda_{1}=\lambda_{2}=1$; and $\lambda_{1}=1$, $\lambda_{2}=0$. Their extrapolated values are $M^{2}_{s}(\infty)=0.0061(7)$, $0.0944(3)$, and $0.0609(2)$ respectively.
In the vertical path with $\lambda_{1}$ fixed at $0.5$ and varying $\lambda_{2}$, two phase transitions occur: one from the hexagon phase to the AFM phase, and the other from the AFM phase to the LSD phase. Through finite-size extrapolation, the critical points are accurately determined to be $\lambda_{2,c}=0.323(7)$ and $2.040(1)$. In the vertical path with $\lambda_{1}$ fixed at $1$ and varying $\lambda_{2}$, only one quantum phase transition is observed, namely from the AFM phase to the LSD phase. The critical point for this transition is accurately identified as $\lambda_{2,c}=2.320(2)$. At the point where $\lambda_{1}=\lambda_{2}=1$, the system is topologically equivalent to the square-lattice Heisenberg model. In the thermodynamic limit, the staggered magnetic order is extrapolated to approximately $0.307$ as shown in Fig. 7, which aligns closely with previously reported results [40]. When $\lambda_{1}=1$ and $\lambda_{2}=0$, the magnetic order is reduced compared to the case where $\lambda_{1}=\lambda_{2}=1$. In the vertical path with $\lambda_{1}$ fixed at $4$ and varying $\lambda_{2}$, the AFM phase region shrinks compared to cases with smaller $\lambda_{1}$ values. Two continuous quantum phase transitions occur between the AFM phase and the OSD phase, as well as between the AFM phase and the LSD phase. The critical points for these transitions are accurately determined to be $\lambda_{2,c}=3.939(6)$ and $4.228(4)$. And the AFM order at the $\lambda_{1}=\lambda_{2}=4$ point is very small compared to the other AFM region, as shown in Fig. 7. To illustrate the quantum phase transition between the Haldane-like SPT phase and the AFM phase, we consider a horizontal path where $\lambda_{1}$ is varied while $\lambda_{2}$ is fixed at $1.4$. As shown in Fig. 6(d), the Binder cumulants for different system sizes exhibit intersections that remain relatively unchanged with increasing system sizes. By performing extrapolation, we accurately determine a quantum critical point at $\lambda_{1,c}=0.106(1)$. This is analogous to the case of coupled Haldane chains transitioning into a 2D square lattice, where a similarly small critical interchain interaction has been reported [41].
We plot the identified phase transition points on the phase diagram and use the same finite-size scaling method to determine additional points across the phase diagram. This process results in a detailed and comprehensive phase diagram for the Heisenberg model on the 2D SHO lattice, as displayed in Fig. 2. Due to the presence of an even sublattice, the AFM phase tends to emerge readily. It occupies a middle area of the phase diagram and is surrounded by four other nonmagnetic phases.
Figure 8: (a-d) The data collapse of the Binder cumulant $U_{2}$ for the phase transitions between the AFM phase and the hexagon phase ($\lambda_{1}=0.5$), LSD phase ($\lambda_{1}=1.0$), OSD phase ($\lambda_{1}=2.0$), and Haldane-like SPT phase ($\lambda_{2}=1.4$), respectively. Here, $\lambda_{1/2,c}$ and $\nu$ represent the optimal critical point and critical exponent, respectively. The error estimation is based on a 95% confidence interval of Gaussian distribution.
@@ III.4 Critical behaviors
After studying the position of each phase in the model clearly, we continue to explore the critical behavior in the phase transition. Since the antiferromagnetic phase breaks the spin $SU(2)$ continuous symmetry in this model while the other phases such as the dimer phase or hexagon phase do not break this continuous symmetry and translation symmetry, it is reasonable to conjecture that the phase transition of this model should belong to the three-dimensional (3D) $O(3)$ universality class. At present, there are precise critical exponent values for the $O(3)$ universality class in the literature [42]. We only need to do the data collapse of the Binder cumulant $U_{2}$ and compare the obtained critical exponent with the known value for confirmation. The finite-size scaling form of the Binder cumulant is given by $U_{2}(\lambda_{1/2},L)=f[L^{1/\nu}(\lambda_{1/2}-\lambda_{1/2,c})/\lambda_{1/2,c}]$, where $f$ is the scaling function, $\lambda_{1/2,c}$ is the critical point, and $\nu$ is the critical exponent of correlation length. To extract the optimal critical point and critical exponent from the finite-size Binder cumulant, we employ the scaling analysis method based on Gaussian process regression proposed in Refs. [43, 44], which enables data from different system sizes to collapse onto a single smooth scaling function at the optimal best-fit critical parameters. We select representative paths from the four phase boundaries in the phase diagram and apply the above scaling analysis method to determine their critical exponents and judge all phase transition types accordingly.
@@
The results of data collapse are shown in Fig. 8, where the best estimates of the critical point $\lambda_{1/2,c}$ and critical exponent $\nu$ for different phase transitions are presented. The critical points obtained here are all consistent with the phase transition points shown in Fig. 2. The critical exponents $\nu$ we obtain here are all in good agreement with the $O(3)$ value $\nu\approx 0.7112$ [42], indicating that the phase transition from the four nonmagnetic phases to the AFM phase in the Heisenberg model on the SHO lattice indeed belongs to the 3D $O(3)$ universality class.
III.5 Stability of the Haldane-like SPT phase
Our previous study of the $J_{1}\text{-}J_{3}$ ladder revealed a Haldane-like SPT phase, intervening between the hexagon and the LSD phases. This phase was identified through characteristic signatures in both the edge spectrum and the entanglement spectrum.
While this model currently exists only in theoretical calculations, its potential realization in magnetic materials highlights the importance of the nearest-neighbor inter-unit-cell antiferromagnetic exchange interaction, $J_{2}$ [denoted by the red lines in the inset of Fig. 9(a)], which cannot be neglected. This leads us to investigate the effect of $J_{2}$ on the SPT phase. Based on the phase diagram obtained from QMC-calculated spin stiffness for a system size of $L=24$, shown in Figs. 9(a) and 9(b), we find that $J_{2}$ does not destroy the SPT phase. On the contrary, it enlarges the region of the SPT phase—a promising result for material realization. To further confirm the SPT phase at $J_{2}=J_{3}=1.4$, we also present the triplet gaps under PBCs and OBCs, as well as the entanglement spectrum, in Figs. 9(c) and 9(d). The emergence of end states and a twofold degenerate entanglement spectrum confirms the presence of the SPT phase.
Figure 9: (a) Spin stiffness along the $x$ direction as a function of $J_{3}$ at fixed $J_{2}$. Peak positions indicate phase boundaries. Inset: Ladder geometry with a rectangular unit cell. (b) $J_{3}\text{-}J_{2}$ phase diagram determined from spin stiffness. A large $J_{2}$ can induce an inter-unit-cell dimer phase, while introducing $J_{2}$ expands the region of the Haldane-like SPT phase. (c) Extrapolation of triplet gaps at $J_{2}=J_{3}=1.4$ under PBCs and OBCs. (d) Entanglement spectrum at $J_{2}=J_{3}=1.4$ for system size $L=32$. (e) Spin stiffness along the $x$ direction as a function of $J_{2}=J_{3}$ at fixed $J_{4}$. (f) Phase diagram of the $J_{3}\text{-}J_{4}$ ladder model with $J_{2}=J_{3}$. Inset: Ladder geometry with a hexagonal structure.
If a material’s lattice is built on a hexagonal unit cell and $J_{3}$ is non-negligible, other third-neighbor interactions within the hexagon—such as $J_{4}$—should also be taken into account. To examine the effect of $J_{4}$, we select a diagonal path with $J_{2}=J_{3}$ in Fig. 9(b) and employ spin stiffness to map out the $J_{4}$ phase diagram, as shown in Figs. 9(e) and 9(f). Our finite-size numerical results indicate that for the SPT phase to be preserved, these third-neighbor interactions $J_{4}$ should generally be weaker than $J_{3}$. Additionally, next-nearest-neighbor interactions within the hexagon constitute another possible perturbation. Although such frustrating interactions may influence the SPT phase, it can remain stable as long as they are sufficiently weak due to the finite topological gap of the SPT phase.
Finally, we note that the preceding discussion on SPT stability is based on the ladder geometry. When extended to two dimensions, our phase diagrams in Fig. 2 indicate that weak inter-ladder couplings ($J_{2}\approx 0.1J_{1}$) can easily destabilize the one-dimensional SPT phase, driving a transition into a magnetically ordered phase.